# EVINCE: Optimizing Adversarial LLM Dialogues via Conditional Statistics and Information Theory

### v1

@@ Abstract
This paper introduces EVINCE (Entropy and Variation IN Conditional Exchanges), a dialogue framework advancing Artificial General Intelligence (AGI) by enhancing versatility, adaptivity, and reasoning in large language models (LLMs). Leveraging adversarial debate and a novel dual entropy theory, EVINCE improves prediction accuracy, robustness, and stability in LLMs by integrating statistical modeling, information theory, and machine learning to balance diverse perspective exploration with strong prior exploitation. The framework’s effectiveness is demonstrated through consistent convergence of information-theoretic metrics, particularly improved mutual information, fostering productive LLM collaboration. We apply EVINCE to healthcare, showing improved disease diagnosis, and discuss its broader implications for decision-making across domains. This work provides theoretical foundations and empirical validation for EVINCE, paving the way for advancements in LLM collaboration and AGI development.
EVINCE: Optimizing Adversarial LLM Dialogues
via Conditional Statistics and Information Theory
Edward Y. Chang Computer Science Clinical Mind AI Stanford University echang@cs.stanford.edu Draft: 2/15/2024; Revised: 8/15/2024
@@ 1 Introduction
The pursuit of Artificial General Intelligence (AGI) remains a central goal of AI research. We propose a paradigm shift in this quest: utilizing multiple Large Language Models (LLMs) engaged in synergistic dialogues as a crucial step towards AGI. This approach, we contend, addresses key limitations of current AI systems and provides a novel pathway to more robust, versatile, and capable artificial intelligence. Specifically, our work targets three core AGI characteristics: versatility, adaptivity, and reasoning capability.
Current LLMs, despite their remarkable capabilities, face significant challenges, including hallucination (generating false or nonsensical information), bias (reflecting and potentially amplifying societal prejudices), and limited reasoning (difficulties in complex problem-solving and logical inference). We posit that multi-agent dialogue systems offer a promising avenue to address these challenges. By fostering diversity and debate among LLMs, these systems can mitigate biases and promote enhanced reasoning capabilities. Furthermore, the iterative nature of multi-round dialogues allows for continuous context enrichment, enabling LLMs to access more precise information and formulate more accurate responses, thus reducing the occurrence of hallucinations.
Our previous work, SocraSynth Chang (2023a), addresses LLM limitations through structured multi-agent dialogues. By leveraging both adversarial and collaborative interactions between LLMs, SocraSynth demonstrates quantifiable improvements across various domains. In healthcare, it achieves a 5% improvement in diagnostic accuracy Chang and et al. (2023), while in news analysis, it identifies and explains biases in articles Chang (2024c). These advancements showcase enhanced versatility, adaptivity, and reasoning capability—key AGI characteristics.
SocraSynth’s adversarial component promotes the exploration of diverse perspectives, while its collaborative component fosters rigorous reasoning to reach well-reasoned conclusions. This synergy has yielded measurable gains beyond healthcare and bias mitigation, extending to geopolitical analysis Chang (2023b), corporate planning Tsao (2023), investment banking Chang (2024b), and emotional behavior modeling Chang (2024a). These results demonstrate SocraSynth’s effectiveness in mitigating LLM limitations and achieving substantial performance improvements across various applications, highlighting its potential for advancing towards AGI’s generalized problem-solving capabilities.
Although SocraSynth has demonstrated effectiveness, it relies on “contentiousness” as a qualitative measure to moderate the linguistic behaviors of participating LLMs. For instance, a high contentiousness value (0.9 out of 1.0) might lead LLMs to challenge each other’s assumptions and propose alternative solutions, while a low value (below 0.3) could encourage them to synthesize their viewpoints and find common ground. While the concept of contentiousness has proven useful in guiding SocraSynth dialogues, its qualitative nature limits its precision and explainability.
In this paper, we extend SocraSynth by introducing three theoretical pillars to quantify “contentiousness” and moderate dialogues based on statistical and information theories. These pillars, collectively referred to as EVINCE (Entropy and Variation IN Conditional Exchanges), provide quantitative measures for justifiable and explainable multi-agent dialogue moderation and evaluation.
These pillars are:
@@
    Inclusiveness Exploration: We develop methods to ensure dialogues explore all potential perspectives. We use conditional statistics via in-context learning Xie et al. (2021) to “free” an LLM agent from its default “maximal likelihood” next-token prediction behavior, allowing it to adopt specific stances. We introduce a dual entropy optimality theory to balance the exploration of new ideas with adherence to priors, thus optimizing information exchange between agents for comprehensive and stable discourse.
@@
    Information Flow Dynamics: We introduce information theory-based metrics to quantify and optimize dialogue dynamics. These measure information diversity (entropy), novelty (statistical divergence scores), and inter-agent persuasion (mutual information). These metrics enable us to assess and enhance the quality and efficiency of information flow within the multi-agent system, fostering rich and productive exchanges.
@@
    Reasoning Quality and Coherence: We establish frameworks to assess the logical structure and coherence of multi-agent reasoning. This pillar evaluates argument validity, analytical depth, and dialogue coherence. We synergistically integrate our previously developed CRIT system Chang (2023c) into SocraSynth. CRIT, which combines Socratic methods with formal reasoning techniques, enhances our ability to conduct critical thinking through evaluating argument quality, information-source credibility, and overall “reasonableness” within the dialogue. This integration ensures that the collective reasoning of LLM agents is not only diverse but also logically sound and aligned with the dialogue’s objectives.
The core strength of SocraSynth and EVINCE in advancing towards AGI lies in their ability to enhance key AGI characteristics through multi-agent dialogues. By employing conditional statistics and information theory, they boost versatility and adaptivity, allowing LLMs to transcend their typical “maximal likelihood” behaviors and mimic how humans adapt their linguistic behaviors to complete tasks. The framework’s debate structure fosters advanced reasoning and critical thinking capabilities. Together, these advancements represent significant steps towards achieving the complex, intricate capabilities required for AGI.
The main contributions of this paper are:
@@
    EVINCE Framework Design: Unlike approaches that use debate merely to improve accuracy via redundancy, EVINCE facilitates information discovery, bias mitigation, and decision-making that requires both breadth and depth of information.
@@
    Theoretical Foundations: We establish a theoretical basis for EVINCE, rooted in conditional Bayesian statistics, mutual information, and dual entropy. These principles are applied to measure, monitor, and modulate collaborative LLM interactions, contributing to a deeper understanding of how LLMs can effectively cooperate for improved decision-making. The dual entropy theory is novel and ground-breaking, illustrating how a productive decision-making process should start with room for diverse input and stable objectives, and then, through information exchange, converge to optimal decision/prediction.
@@
    Empirical Validation: We provide empirical validation of EVINCE’s underlying theories, highlighting the framework’s effectiveness in balancing exploration and exploitation to enhance prediction accuracy. We also introduce a set of maxims derived from our empirical findings, offering practical guidance for optimizing mutual information and minimizing various divergence measures.
The remainder of this paper is structured as follows: Section 2 discusses related work and addresses common misconceptions about our approach. Section 3 presents the three pillars of EVINCE, including its maxims, theories, and algorithms. Section 4 provides empirical evidence of EVINCE’s benefits by revisiting and expanding upon previous SocraSynth experiments in healthcare and bias detection. Section 5 offers concluding remarks and discusses future directions.
2 Related Work
The core objective of adversarial debate, as embodied in SocraSynth with EVINCE, is to foster diverse opinions and challenge assumptions, ultimately leading to more comprehensive and informed decision-making. This contrasts with traditional ensemble learning methods, which prioritize error diversity for improved accuracy. This section surveys ensemble learning and adversarial learning, highlighting the unique aspects of EVINCE and its theoretical underpinnings.
2.1 Ensemble Learning
Ensemble methods, such as Bagging Breiman (1996), Boosting Freund and Schapire (1997), and Mixtures of Experts Jacobs et al. (1991), traditionally aim to enhance machine learning performance by combining predictions from multiple models to achieve error diversity. Similarly, early LLM debate frameworks, e.g., Michael et al. (2023); Chan et al. (2023); Liang et al. (2023); Du et al. (2023) have focused primarily on improving answer accuracy rather than fostering a comprehensive exploration of information for critical decision-making.
SocraSynth distinguishes itself from traditional ensemble methods by prioritizing the generation of diverse predictions over the mere avoidance of errors. This is achieved through a dynamic protocol that adaptively adjusts the “contentiousness” level of the debate, encouraging models to initially explore a wide range of perspectives and rigorously assess the quality of arguments. Through constructive exchange of ideas in subsequent iterations, SocraSynth converges on a more informed and robust decision. Its approach, grounded in EVINCE’s dual entropy theory, allows for the exploration of various ideas by one LLM while maintaining stability through the other, effectively preventing premature convergence on a narrow set of perspectives.
2.2 Addressing Common Misconceptions about Adversarial LLM Dialogue
The rapid growth in academic conference submissions has led to an expanded reviewer pool, potentially impacting review quality. In this context, some misconceptions about our work have arisen, which we address here:
  1. 1.
    Misconception: Multiple LLMs cannot mitigate hallucination.
    Some reviewers have expressed skepticism about the ability of multiple LLMs to reduce hallucination. However, our approach does not simply aggregate potentially flawed outputs. Instead, SocraSynth leverages the strengths of multiple models to cross-check, challenge, and refine information. This process is analogous to peer review in scientific discourse, where collective scrutiny enhances the reliability of conclusions.
  2. 2.
    Misconception: Multiple LLM dialogues cannot lead to knowledge discovery.
    Despite evidence provided in Chang (2023a) and Chang and Chang (2023), some reviewers remain skeptical about the potential for knowledge discovery through LLM dialogues. Our work demonstrates that carefully structured interactions between LLMs can indeed lead to novel insights and knowledge synthesis by leveraging the polydisciplinary knowledge representation inherent in LLMs.
  3. 3.
    Misconception: The approach lacks theoretical foundation.
    Our approach is grounded in conditional statistics and information theory, which we use to optimize LLM interactions. This theoretical foundation ensures that the dialogue converges towards more accurate and less biased outcomes, addressing concerns about the rigor of our methodology.
By addressing these misconceptions, we aim to clarify the scientific basis and potential of our approach. The synergistic interaction between LLMs in SocraSynth, coupled with EVINCE’s theoretical pillars, represents a novel method for enhancing the capabilities of language models. This approach has applications in reducing bias, mitigating hallucination, and potentially advancing the field towards more robust AI systems.
Metric Pros Cons Remedies
Cross Entropy (CE) Shore and Johnson (1980) Measures how well the predictions of one model fit the actual distribution of another model’s outputs (asymmetric). Computationally intensive especially with large models and data sets; sensitive to the exact nature of probability distributions. Optimize computation strategies; use approximations or sampling methods to manage large data sets or complex models.
Entropy Shannon
Shannon (1948)
Indicates level of diversity; high suggests exploration of possibilities, and low for confidence on few choices High entropy might indicate noise rather than useful diversity; low entropy might mask important variability. Use critical reading methods (Appendix A) to assess argument quality; implement noise detection to differentiate between useful diversity and noise.
Jensen-Shannon
Divergence (JS)
Lin (1991)
KL Divergence
Kullback (1951)
Mutual Info (MI) Shore and Johnson (1980) Measures reduction of uncertainty; symmetric. Does not indicate the directionality of information flow. Supplement with directional information metrics; normalized with max entropy of A and B.
Wasserstein
Distance (WD)
Kantorovich (1942)
Direct measure of how similar or different the model outputs are; it depicts symmetric relationship. Not bounded but can be normalized or bounded for consistent interpretation. Define context-specific bounds for low, medium, and high divergence; consider normalizing it for non-directional comparisons.
Table 1: Summary of metrics for assessing LLM debates (equations are presented in Appendix D)
3 Algorithm, Maxims, and Theories
@@
Organize a structured debate between two equally competent large language models (LLMs), LLMA and LLMB, to conduct $t$ rounds. At each round $t$, each model produces a probability distribution, denoted as $P_{A}^{(t)}$ and $P_{B}^{(t)}$, over $C$ possible outcomes, accompanied by supporting arguments $R_{A}^{(t)}$ and $R_{B}^{(t)}$. The goal is to design an iterative debate process that leverages the structured exchange of arguments to enable the models to converge on an optimal prediction distribution $P^{*}$ across the $C$ classes. Optimality is defined as achieving the highest possible accuracy on the prediction task.
To address this problem, we first introduce key metrics from information theory that form the basis of our approach.
@@ 3.1 Preliminaries in Information Theory
This section presents the key metrics used to measure information diversity, similarity, divergence, and other relevant factors within EVINCE. These metrics serve three primary objectives:
3.1.1 Fostering diversity of perspectives while ensuring reasoning quality
  • Wasserstein Distance (WD): Measures distribution difference between predictions to identify exploration opportunities Kantorovich (1942); Rubner et al. (2000); Villani (2008).
  • Shannon Entropy or Relative Entropy: Measures diversity of perspectives Cover and Thomas (2006); Shannon (1948).
  • Reasoning Quality: The CRIT algorithm evaluates the logical soundness and persuasiveness of supporting arguments Chang (2023c), helping to identify and mitigate hallucinations and poorly-reasoned arguments.
3.1.2 Exploring new possibilities while adhering to the dialogue subject
  • Correlation Coefficients: Tracks the evolution of opinions and assesses debate stability Brown et al. (2005) toward the goal of the dialogue.
  • Mutual Information (MI): Quantifies information overlap to ensure focused and productive debates Cover and Thomas (2006) and measures the degree of agreement/disagreement.
3.1.3 Examining information convergence and establishing termination criteria
  • Jensen-Shannon (JS) Divergence: Assesses similarity between probability distributions Lin (1991) (symmetric).
  • Cross Entropy (CE): Measures the asymmetric difference between prediction distributions Shore and Johnson (1980).
  • Kullback-Leibler (KL) Divergence: Reveals asymmetric differences between probability distributions Kullback (1951).
Table 1 summarizes these metrics and their pros and cons. Next, we present the EVINCE algorithm and explain how these metrics are used to moderate an LLM dialogue to achieve a balance between exploration and exploitation, leading to optimal prediction outcomes. In Section 3.3, we will introduce several maxims and prove the dual entropy theorem, which provides a theoretical foundation for the EVINCE algorithm.
@@ 3.1.3 Examining information convergence and establishing termination criteria
Figure 1: Specifications of Algorithm EVINCE. Key points: 1) Asymmetric Start: In Step #1, LLMA initiates with opening arguments based solely on the given information, while LLMB starts with access to LLMA’s prediction and arguments for refutation. 2) Termination Criteria: The while loop in Step #2 considers three factors: Wasserstein distance, mutual information, and argument quality. EVINCE terminates if the dialogue ceases to make significant progress. 3) Further Details: Maxims #1 to #4 provide additional explanations of the algorithm’s principles. 4) Argument Evaluation: For details on the CRIT algorithm used to assess argument quality, refer to Chang (2023c).
3.2 EVINCE Algorithm Specifications
Figure 1 formally specifies the detailed operations of EVINCE.
Setup and Initialization:
EVINCE employs two equally competent LLM instances, $LLM_{A}$ and $LLM_{B}$, which can be different models (e.g., GPT and Claude) or independent instances of the same model. Given an information set $S$ and a class-label set $C$, EVINCE outputs a probability distribution over $C$ with justifications. For example, $S$ could represent a patient’s symptoms and $C$ a set of diseases, with EVINCE moderating a dialogue to predict the patient’s disease(s).
The debate “contentiousness” is initially set to 90% to encourage disagreement and explore diverse perspectives. This high level is maintained for the first two iterations, then gradually reduced by $x\%$ in each subsequent iteration to shift from exploration to exploitation [^1] The setting of $x$ at the beginning of a dialogue iteration can depend on the values of all moderation subroutines. Our empirical experience suggests $x=25\%$ because a typical dialogue concludes in five iterations, and $x=20\%$ would set contentiousness to be 90%, 90%, 68%, 50%, and 38%.. Table 2 demonstrates how GPT adapts its linguistic behavior based on the “contentiousness” parameter, showcasing the model’s ability to adjust to different contextual settings. Importantly, this table is generated by GPT itself, not by human moderators, as detailed in Chang (2023a) and further explored in Section 4.
C.L. Tone Emphasis Language
$0.9$ Highly confrontational; focused on raising strong ethical, scientific, and social objections. Flagging risks and downsides; ethical quandaries, unintended consequences, and exacerbation of inequalities. Definitive and polarizing, e.g., “should NOT be allowed,” “unacceptable risks,” “inevitable disparities.”
$0.7$ Still confrontational but more open to potential benefits, albeit overshadowed by negatives. Acknowledging that some frameworks could make it safer or more equitable, while cautioning against its use. Less polarizing; “serious concerns remain,” “needs more scrutiny.”
0.5 Balanced; neither advocating strongly for nor against gene editing. Equal weight on pros and cons; looking for a middle ground. Neutral; “should be carefully considered,” “both benefits and risks.”
0.3 More agreeable than confrontational, but maintaining reservations. Supportive but cautious; focus on ensuring ethical and equitable use. Positive but careful; “transformative potential,” “impetus to ensure.”
0.0 Completely agreeable and supportive. Fully focused on immense potential benefits; advocating for proactive adoption. Very positive; “groundbreaking advance,” “new era of possibilities.”
Table 2: Changes in linguistic behaviors of LLMs at different contentiousness levels.
@@ C.L. Tone Emphasis Language
EVINCE employs four key subroutines to manage the exploration-exploitation tradeoff, ensuring information diversity, quality, and stability:
  • CRIT: Evaluates argument quality and source credibility. Low scores can trigger dialogue termination.
@@
  • Additional metrics: KL divergence, Jensen-Shannon divergence, and cross entropy ensure evaluation consistency (see Table 1).
When all metrics reach a plateau, indicating no further improvement, EVINCE terminates the dialogue. At this point, both LLM instances are prompted to jointly deliver a conciliatory conclusion with comprehensive arguments and counterarguments. Additionally, they provide a list of missing information that, if obtained, could potentially improve prediction accuracy and reliability.
@@
Given the contentiousness level set for a dialogue iteration, each LLM generates a new prediction distribution on $C$ with arguments, based on the other’s last prediction and accumulated reasons as the context.
@@
  • Weight predictions based on supporting argument quality (evaluated by CRIT).
  • Handle uncertainty by soliciting missing information when final prediction entropy is high (see Chang (2024b) for RAG application).
EVINCE creates a structured debate between AI models, balancing diverse viewpoints with consensus-building to produce accurate, well-reasoned predictions.

### v4

# EVINCE: Optimizing Multi-LLM Dialogues Using Conditional Statistics and Information Theory

@@ Abstract
$\mathsf{EVINCE}$ (Entropy and Variation IN Conditional Exchanges) is a novel framework for optimizing multi-LLM dialogues using conditional statistics and information theory. It addresses limitations in multi-agent debate (MAS) frameworks, where multiple LLMs “chat” without behavior modulation or mutual information quality assessment. Using dual entropy optimization to balance perspective diversity and prior knowledge, $\mathsf{EVINCE}$ provides quantitative tools to dynamically regulate LLM linguistic behaviors. When mutual information is low and both cross-entropy and Wasserstein distance are high, $\mathsf{EVINCE}$ promotes contentious dialogues to expose diverse perspectives and uncover inconsistencies. Conversely, as cross-entropy decreases and mutual information stabilizes, it transitions discussions into a conciliatory phase, encouraging compromise and acknowledgment of valid points. Using information-theoretic metrics and optimizing mutual information, $\mathsf{EVINCE}$ emerges as a structured and highly effective framework for multi-LLM collaboration.
Machine Learning, Multiple Agents
@@ 1 Introduction
Current large language models (LLMs) (e.g. (Anthropic, 2024; OpenAI, 2023; Gemini, 2023)) demonstrate remarkable capabilities but face significant challenges, including hallucination (generating false information), bias (reflecting societal prejudices), and limited reasoning (difficulties in complex problem solving and logical inference). These challenges critically limit LLMs’ reliability and applicability in real-world scenarios.
Multi-agent dialogue systems present a promising approach to addressing these limitations. By fostering diversity and structured debate among LLMs, these systems can mitigate biases and enhance reasoning through collaborative intelligence.
However, most prior works in Multi-Agent Debate (MAD) (e.g., (Abdelnabi et al., 2024; Chan et al., 2023; Fu et al., 2023; Li et al., 2023; Liang et al., 2023; Michael et al., 2023; Smit et al., 2024)) primarily follow the traditional Mixture of Experts (MoE) paradigm (Jacobs et al., 1991). These methods rely on redundancy to mask errors (Freund & Schapire, 1997) and leverage error diversity to improve response quality. As a result, their debates often suffer from repetitive arguments, reinforcement of flawed reasoning, and inconsistent conclusions (Wang et al., 2024; Zhang et al., 2024). Moreover, these systems lack dynamic modulation of contentious intensity during debates, leading to discussions that are either uniformly mild or overly contentious, ultimately producing suboptimal results.
Recent work by SocraSynth (Chang, 2023a) explored the dynamic modulation of debate “contentiousness” to generate diverse perspectives and then refine with reasoning to improve results. However, its qualitative approach to modulating contentiousness in linguistic behavior limits precise control and optimization of these interactions.
In this paper, we advance MAD by introducing $\mathsf{EVINCE}$ (Entropy and Variation IN Conditional Exchanges), which provides quantitative foundations for multi-agent dialogue moderation. $\mathsf{EVINCE}$ builds on three theoretical pillars:
@@
    Inclusive Exploration: We develop methods to ensure dialogues comprehensively explore diverse perspectives. Using conditional statistics, we adjust an LLM agent’s behavior beyond its default “maximum likelihood” next-token prediction, enabling it to adopt specific stances via in-context learning. To balance idea exploration with adherence to prior knowledge, we introduce a dual-entropy optimization framework, improving information exchange for richer discourse.
@@
    Information Flow Dynamics: We quantify and optimize dialogue interactions using information-theoretic metrics, measuring information diversity (entropy), novelty (divergence scores), and inter-agent persuasion (mutual information). These metrics enhance the quality and efficiency of information flow in multi-agent settings, fostering richer, more productive exchanges.
@@
    Reasoning Quality and Coherence: We establish frameworks to evaluate the logical structure and coherence of multi-agent reasoning. This includes assessing argument validity, analytical depth, and dialogue consistency. We integrate the CRIT algorithm (Chang, 2023b), which combines Socratic methods with formal reasoning, to enhance argument evaluation and ensure logically sound, goal-oriented discourse.
The contributions of this paper are as follows:
@@
    Framework Design: Unlike prior debate-based methods that enhance accuracy primarily through redundancy, $\mathsf{EVINCE}$ enables structured information discovery, bias mitigation, and decision-making by integrating both broad exploratory analysis and deep evaluative reasoning.
@@
    Theoretical Foundations: $\mathsf{EVINCE}$ establishes a theoretical foundation rooted in Bayesian statistics, mutual information, and dual entropy. The dual entropy framework formalizes how decision-making should begin with diverse, high-entropy inputs and stable objectives, then iteratively refine through information exchange to achieve optimal predictions.
@@
    Empirical Validation: Through experiments in disease diagnosis and news debiasing (Section 3), we validate $\mathsf{EVINCE}$’s principles and demonstrate how optimizing information flow improves predictive accuracy and reasoning robustness.
2 EVINCE Algorithm, Maxims, and Theories
@@
Organize a structured debate between two equally competent large language models (LLMs), LLMA and LLMB, to conduct $t$ rounds. At each round $t$, each model generates confidence scores for its top-k predictions, denoted as $P_{A}^{(t)}$ and $P_{B}^{(t)}$, over $C$ possible outcomes, accompanied by supporting arguments $R_{A}^{(t)}$ and $R_{B}^{(t)}$. The goal is to design an iterative debate process that leverages the structured exchange of arguments to enable the models to converge on an optimal prediction ranking $P^{*}$ across the $C$ classes. Optimality is defined as achieving the highest possible accuracy on the prediction task while maintaining strong reasoning support.
@@ 2.1 Preliminaries in Information Theory
This section summarizes the key metrics used in $\mathsf{EVINCE}$ to evaluate information diversity, similarity, divergence, and other relevant factors. A detailed comparison, including advantages and limitations, is provided in Appendix A. Notably, computational complexity is not a concern in our context, as we work with probability distributions of up to ten categories. These metrics serve three primary objectives:
A. Fostering diversity of perspectives while ensuring reasoning quality:.
    Wasserstein Distance (WD): Measures the difference between prediction distributions to identify exploration opportunities (Kantorovich, 1942; Rubner et al., 2000; Villani, 2008).
    Shannon Entropy or Relative Entropy: Measures diversity of perspectives (Shannon, 1948).
    Reasoning Quality: The $\mathsf{CRIT}$ (Critical Thinking) algorithm evaluates the logical soundness and persuasiveness of supporting arguments (depicted in Appendix B), helping to identify and mitigate hallucinations and poorly-reasoned arguments (Chang, 2023b).
B. Exploring new possibilities while maintaining stability toward the goal of the dialogue:.
    Correlation Coefficients: Tracks the evolution of opinions and assesses debate stability (Brown et al., 2005) toward the goal of the dialogue.
    Mutual Information (MI): Quantifies information overlap to ensure focused and productive debates (Cover & Thomas, 2006b) and measures the degree of agreement/disagreement.
C. Examining information convergence and establishing dialogue termination criteria:.
    Jensen-Shannon (JS) Divergence: Assesses similarity between probability distributions (symmetric).
    Cross Entropy (CE): Measures the asymmetric difference between prediction distributions.
    Kullback-Leibler (KL) Divergence: Reveals asymmetric differences between probability distributions.
Next, we present the $\mathsf{EVINCE}$ algorithm and explain how these metrics are used to moderate an LLM dialogue to achieve a balance between exploration and exploitation, leading to optimal prediction outcomes.
2.2 Algorithm Specifications
Figure 1 specifies the detailed algorithm of $\mathsf{EVINCE}$. It employs two equally competent LLM instances, $LLM_{A}$ and $LLM_{B}$, which can be different models (e.g., GPT and Claude) or independent instances of the same model. Given an information set $S$ and a class-label set $C$, $\mathsf{EVINCE}$ outputs a confidence distribution over $C$ with justifications. For example, $S$ could represent a patient’s symptoms and $C$ a set of diseases, with $\mathsf{EVINCE}$ moderating a dialogue to predict the patient’s disease(s).
@@ 2.2 Algorithm Specifications
Figure 1: Specifications of Algorithm $\mathsf{EVINCE}$. Key points: 1) Asymmetric Start: In Step #1, LLMA initiates with opening arguments based solely on the given information, while LLMB starts with access to LLMA’s prediction and arguments for refutation. 2) Termination Criteria: The while loop in Step #2 considers three factors: Wasserstein distance, mutual information, and argument quality. $\mathsf{EVINCE}$ terminates if the dialogue ceases to make significant progress. 3) Further Details: Maxims #1 to #4 provide additional explanations of the algorithm’s principles. 4) Argument Evaluation: Step #2.2 evaluates of argument quality, and the while loop examines if argument quality continues to improve. 5) Update($\Delta$) modulates contentiousness, and see Maxim #3 for its specifications.
@@
$\mathsf{EVINCE}$ employs four key subroutines from the previous subsection to manage the exploration-exploitation tradeoff, ensuring information diversity, quality, and stability:
  • $\mathsf{CRIT}$: Evaluates argument quality and source credibility. Low scores can trigger dialogue termination.
@@
  • Additional metrics: KL Jensen-Shannon divergence, and cross entropy ensure evaluation consistency.
When all metrics plateau, indicating no further improvement, $\mathsf{EVINCE}$ terminates the dialogue. (See the while loop in the second step of the algorithm).) At this stage, both LLM instances are prompted to collaboratively deliver a conciliatory conclusion that includes comprehensive arguments and counterarguments. They also provide a list of missing information that, if obtained, could enhance prediction accuracy and reliability, potentially using Retrieval-Augmented Generation (RAG) techniques.
@@
Given the contentiousness level set for a dialogue iteration, each LLM generates a new prediction distribution on $C$ with supporting arguments, considering the other LLM’s previous prediction and accumulated reasoning as context. The contentiousness level updates follow Maxims #2.2 and #3. The top-k predictions from LLMs are then calibrated using reasoning quality assessment following Maxim #4.
@@
  • Weight predictions based on supporting argument quality (evaluated by $\mathsf{CRIT}$).
  • Handle uncertainty by soliciting missing information when final prediction entropy is high (Maxim #4.4).
$\mathsf{EVINCE}$ facilitates a structured debate between AI models, unearthing all perspectives related to the prediction tasks at hand while balancing diverse viewpoints and fostering consensus to produce accurate, well-reasoned predictions.